Evaluate ∫∫Rcos(x^2+y^2)dA, where R is the region above the x-axis within the circle x^2+y^2= 25.
1
Expert's answer
2020-11-05T10:33:18-0500

"\displaystyle

\textsf{We convert to polar co-ordinates}\\

\mathrm{d}A = r\mathrm{d}r\mathrm{d}\theta\\

x = r\cos\theta\\

y = r\sin\theta\\

x^2 + y^2 = r^2\\

\textsf{Since}\, R\, \textsf{is the region above the}\, x-\textsf{axis},\\\textsf{we only consider the first and}\\\textsf{second quadrants, and thus evaluate}\\\textsf{the double integral.}\\

\theta\, \textsf{is from}\, 0 \, \textsf{to}\, \pi\\

r\, \textsf{is from}\, 0 \, \textsf{to}\, 5\\

\begin{aligned}
\therefore \iint_R \cos(x^2 + y^2) \, \mathrm{d}A &= \int_0^5\int_0^\pi \cos(x^2 + y^2) \, r\mathrm{d}\theta \mathrm{d}r\\
&= \pi\int_0^5 \cos(r^2) \, r \mathrm{d}r\\
&= \frac{\pi\sin(r^2)}{2} \biggr\vert_0^5\\
&= \frac{\pi\sin(25)}{2}
\end{aligned}\\

\textsf{The area of the region is}\,\frac{\pi\sin(25)}{2}"